# Multilingual Performance of a Multimodal Artificial Intelligence System on Multisubject Physics Concept Inventories

### v1

@@ IV.1 Scores by inventory and language
A second key point is that some inventories appear to have stable performance across multiple languages, while others do not. For example, QMCS shows high and relatively stable performance (78–86%) in languages for which data is available, indicating that the AI handles that conceptual domain and its translations fairly well. In contrast, MBT scores fluctuate more dramatically (e.g., 27% in Persian, 53% in Finnish, and 44% in Italian and Hungarian), indicating that performance may depend not just on the conceptual domain but also on the subtleties of the language and phrasing used.
One such outlier is the TUG-K2.6, where performance differs widely across the few languages tested, from 48% in Arabic to 67% in Finnish, with little intermediate data. Similarly, TUG-K3.0-4.0 shows a scattered pattern: 44% in Greek, 55% in English, 60% in Spanish, and dips to 45% or lower in other languages, suggesting highly variable understanding.
| AR | BN | CA | CS | DE | EL | EN | ES | FA | FI | FIL | FR | HE | HI | HR | HU | ID | IS | IT | JA | MS | NL | NO | PA | PL | PT | RU | SK | SR | SV | TA | TH | TR | UK | ZH | |
| ADT | - | - | - | - | - | - | 78 | 81 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 83 | - | - | - | - | - |
| BEMA | - | - | - | - | - | - | 66 | 49 | - | - | - | - | - | - | - | - | - | - | - | 53 | - | - | - | - | - | 57 | - | - | - | 51 | - | - | - | - | 44 |
| CCI | - | - | - | 79 | - | - | 83 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| CDPA | - | - | - | - | - | - | 33 | 40 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| CSEM | - | - | - | - | - | - | 54 | 53 | - | - | - | - | - | - | - | - | 38 | - | - | - | 53 | - | - | - | - | - | - | - | - | 45 | - | - | - | - | - |
| CTSR | - | - | - | - | - | - | 86 | 74 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 62 | 76 | - | - | - | - | - |
| DIRECT | - | - | - | - | 45 | 53 | 49 | 33 | - | 52 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 49 | - | - | - | - | 59 |
| DS | - | - | - | - | 79 | - | 97 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| ECA | - | - | - | - | - | - | 79 | - | - | - | - | - | - | - | 72 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| ECCE | - | - | - | - | - | - | 61 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| EMCA | - | - | - | - | - | - | 78 | - | - | - | - | - | - | - | - | - | 78 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| EMCS | - | - | - | - | - | - | 79 | - | - | 75 | - | - | - | - | - | - | 75 | - | - | - | - | - | - | - | - | - | - | - | - | 71 | - | - | - | - | - |
| FCI | 58 | 39 | 67 | 62 | 68 | 66 | 69 | 68 | 52 | 67 | 59 | 70 | 50 | 49 | 67 | 67 | - | 66 | 72 | 60 | 57 | 62 | 69 | 20 | 73 | 74 | 56 | 62 | - | 64 | 22 | 47 | 66 | - | 57 |
| FMCE | - | - | - | - | - | - | 72 | 81 | - | - | - | - | - | - | - | - | 74 | - | - | 39 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| FORT | - | - | - | - | - | - | 70 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| FTGOT | - | - | - | - | - | - | 27 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 23 | - | - |
| FVA | - | - | - | - | - | - | 76 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| GECI | - | - | - | - | - | - | 82 | - | - | - | - | - | - | - | - | - | - | - | - | 82 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| HTCE | - | - | - | - | - | - | 76 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 68 |
| IBCDC | - | - | - | - | - | - | 62 | - | - | - | - | 64 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| IBCM | - | - | - | - | - | - | 74 | - | - | - | - | 66 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| LPCA | - | - | - | - | - | - | 84 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| LPCI | - | - | - | - | - | - | 65 | 62 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| LSCI | - | - | - | - | - | - | 73 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| MBT | - | - | - | - | 41 | 44 | 44 | 40 | 27 | 53 | - | 41 | - | - | - | - | - | - | 38 | 44 | 38 | - | - | - | - | 35 | - | - | - | - | - | - | 37 | - | - |
| MCS | - | - | - | - | - | - | 58 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| MUQ | - | - | - | - | - | - | 37 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| MWCS1 | - | - | - | - | - | - | 79 | 76 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 35 | - | - | - |
| MWCS2 | - | - | - | - | - | - | 68 | 62 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| NGCI | 76 | - | - | - | - | - | 87 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| NGPSD | - | - | - | - | - | - | 95 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| PIQL | - | - | - | - | - | - | 80 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| QMCA5 | - | - | - | - | - | - | 74 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 67 | - | - | - | - | - | - | - | - | - |
| QMCA6 | - | - | - | - | - | - | 81 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| QMCS | - | - | - | - | - | - | 86 | - | - | 81 | - | - | - | - | - | - | - | - | - | 78 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| QMFPS | - | - | - | - | - | - | 60 | 49 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| QMS | - | - | - | - | - | - | 69 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| QMVI | - | - | - | - | - | - | 32 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| QPCS | - | - | - | - | - | - | 71 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 40 | - | - | - |
| RAPT | - | - | - | - | - | - | 81 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| RCI | - | - | - | - | - | - | 75 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| RFCI | - | - | - | - | - | - | 80 | - | - | 80 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| RKI | - | - | - | - | - | - | 70 | - | - | - | - | 67 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| RRMCS | - | - | - | - | - | - | 74 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| SGCE | - | - | - | - | - | - | 60 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| SPCI | - | - | - | - | - | - | 95 | 94 | - | - | - | - | - | - | - | - | - | - | - | 86 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| STPFaSLSh | - | - | - | - | - | - | 86 | - | - | - | - | - | - | - | - | - | 70 | - | - | - | - | - | - | - | - | 72 | - | - | - | - | - | - | - | - | 59 |
| STPFaSLlo | - | - | - | - | - | - | 97 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| TCE | - | - | - | - | - | - | 88 | - | - | - | - | - | - | - | - | - | - | - | - | 85 | - | - | - | - | - | 78 | - | - | - | - | - | - | - | - | 59 |
| TCS | - | - | - | - | - | - | 79 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 59 | - | - | 69 |
| TOAST | - | - | - | - | - | - | 83 | - | - | - | - | - | - | - | - | - | - | - | - | 80 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| TUG-K2.6 | 48 | - | - | - | 60 | - | - | - | - | 60 | - | 67 | 43 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| TUG-K3.0-4.0 | - | - | - | - | - | 44 | 55 | 60 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 65 | - | - | - | 50 | - | - | - | 45 | 45 |
| TUV | 58 | - | - | - | - | - | 60 | 62 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |

### v3

@@ IV.2 How does language influence the performance of the AI system?
Similarly, MBT scores fluctuate widely — e.g., 27% in Persian, 53% in Finnish, and 44% in Italian and Hungarian (SD=; Range=) — with Finnish performance nearly double that of Persian. In contrast, some inventories show more consistent performance across multiple languages. For example, QMCS shows high and relatively stable scores (78–86%) in the available languages (SD=; Range=), indicating that the AI handles this conceptual domain and its translations fairly well. TUV, ADT, and FMCE, also show relatively stable performance in the 60–80% range. Therefore, these results suggest that performance stability across languages is inventory-dependent.
Out of 36 inventories available in both English and at least one other language, English was the best-performing language in 27 cases (75%). Figure 4 shows the relative performance by language compared to English. For some languages, only a single inventory was available — usually the FCI. Notably for Finnish and French, where we have several translated inventories available, the performance is similar to that in English, with a number of other languages, mostly European, trailing closely behind. On the other hand, performance is much lower in languages such as Persian and Thai.
Merely examining performance on an inventory in a given language is insufficient to determine whether the AI struggles with the same items across languages. Although a detailed item-level analysis is beyond the scope of this manuscript, we provide a brief investigation to assess whether items that the AI finds challenging in English remain so in their non-English versions. One indicator of item difficulty is the agreement among the three independent responses and how consistently the AI system provides the correct answers. Since all inventories — except TUG-K2.6 — were presented to the AI in English, we can directly compare GPT-4o’s performance on English items to their translated counterparts. Our findings indicate that items deemed difficult in English tend to remain challenging in other languages, and vice versa. When all three English responses were incorrect, the correct response rate for the non-English versions was only 18%. This increased to 33% when one English response was correct, 48% for two, and 82% for three. Table 9 in Appendix C illustrates that this trend — where the same items are consistently challenging across various languages — holds true for most languages in our dataset.
An interesting finding is that GPT-4o often exhibits language-switching behavior with non-English inventory items. Portuguese and Spanish were the only two languages where the majority of answers (56% and 59%, respectively) were entirely in the nominal language. In all other cases, the model predominantly switched into English, either fully or in part. Table 8 in Appendix C provides additional details on this language-switching behavior.
Arabic |
Bengali |
Catalan |
Chinese |
Croatian |
Czech |
Dutch |
English |
Filipino |
Finnish |
French |
German |
Greek |
Hebrew |
Hindi |
Hungarian |
Icelandic |
Indonesian |
Italian |
Japanese |
Malay |
Norwegian |
Persian |
Polish |
Portuguese |
Punjabi |
Russian |
Serbian |
Slovak |
Spanish |
Swedish |
Tamil |
Thai |
Turkish |
Ukrainian |
| ADT | - | - | - | - | - | - | - | 78 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 81 | 83 | - | - | - | - |
| BEMA | - | - | - | 44 | - | - | - | 66 | - | - | - | - | - | - | - | - | - | - | - | 53 | - | - | - | - | 57 | - | - | - | - | 49 | 51 | - | - | - | - |
| CCI | - | - | - | - | - | 79 | - | 83 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| CDPA | - | - | - | - | - | - | - | 33 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 40 | - | - | - | - | - |
| CSEM | - | - | - | - | - | - | - | 54 | - | - | - | - | - | - | - | - | - | 38 | - | - | 53 | - | - | - | - | - | - | - | - | 53 | 45 | - | - | - | - |
| CTSR | - | - | - | - | - | - | - | 86 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 62 | - | 74 | 76 | - | - | - | - |
| DIRECT | - | - | - | 59 | - | - | - | 49 | - | 52 | - | 45 | 53 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 33 | 49 | - | - | - | - |
| DS | - | - | - | - | - | - | - | 97 | - | - | - | 79 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| ECA | - | - | - | - | 72 | - | - | 79 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| ECCE | - | - | - | - | - | - | - | 61 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| EMCA | - | - | - | - | - | - | - | 78 | - | - | - | - | - | - | - | - | - | 78 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| EMCS | - | - | - | - | - | - | - | 79 | - | 75 | - | - | - | - | - | - | - | 75 | - | - | - | - | - | - | - | - | - | - | - | - | 71 | - | - | - | - |
| FCI | 58 | 39 | 67 | 57 | 67 | 62 | 62 | 69 | 59 | 67 | 70 | 68 | 66 | 50 | 49 | 67 | 66 | - | 72 | 60 | 57 | 69 | 52 | 73 | 74 | 20 | 56 | - | 62 | 68 | 64 | 22 | 47 | 66 | - |
| FMCE | - | - | - | - | - | - | - | 72 | - | - | - | - | - | - | - | - | - | 74 | - | 39 | - | - | - | - | - | - | - | - | - | 81 | - | - | - | - | - |
| FORT | - | - | - | - | - | - | - | 70 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| FTGOT | - | - | - | - | - | - | - | 27 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 23 | - |
| FVA | - | - | - | - | - | - | - | 76 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| GECI | - | - | - | - | - | - | - | 82 | - | - | - | - | - | - | - | - | - | - | - | 82 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| HTCE | - | - | - | 68 | - | - | - | 76 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| IBCDC | - | - | - | - | - | - | - | 62 | - | - | 64 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| IBCM | - | - | - | - | - | - | - | 74 | - | - | 66 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| LPCA | - | - | - | - | - | - | - | 84 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| LPCI | - | - | - | - | - | - | - | 65 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 62 | - | - | - | - | - |
| LSCI | - | - | - | - | - | - | - | 73 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| MBT | - | - | - | - | - | - | - | 44 | - | 53 | 41 | 41 | 44 | - | - | - | - | - | 38 | 44 | 38 | - | 27 | - | 35 | - | - | - | - | 40 | - | - | - | 37 | - |
| MCS | - | - | - | - | - | - | - | 58 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| MUQ | - | - | - | - | - | - | - | 37 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| MWCS1 | - | - | - | - | - | - | - | 79 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 76 | - | - | 35 | - | - |
| MWCS2 | - | - | - | - | - | - | - | 68 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 62 | - | - | - | - | - |
| NGCI | 76 | - | - | - | - | - | - | 87 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| NGPSD | - | - | - | - | - | - | - | 95 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| PIQL | - | - | - | - | - | - | - | 80 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| QMCA5 | - | - | - | - | - | - | - | 74 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 67 | - | - | - | - | - | - | - | - | - | - |
| QMCA6 | - | - | - | - | - | - | - | 81 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| QMCS | - | - | - | - | - | - | - | 86 | - | 81 | - | - | - | - | - | - | - | - | - | 78 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| QMFPS | - | - | - | - | - | - | - | 60 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 49 | - | - | - | - | - |
| QMS | - | - | - | - | - | - | - | 69 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| QMVI | - | - | - | - | - | - | - | 32 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| QPCS | - | - | - | - | - | - | - | 71 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 40 | - | - |
| RAPT | - | - | - | - | - | - | - | 81 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| RCI | - | - | - | - | - | - | - | 75 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| RFCI | - | - | - | - | - | - | - | 80 | - | 80 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| RKI | - | - | - | - | - | - | - | 70 | - | - | 67 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| RRMCS | - | - | - | - | - | - | - | 74 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| SGCE | - | - | - | - | - | - | - | 60 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| SPCI | - | - | - | - | - | - | - | 95 | - | - | - | - | - | - | - | - | - | - | - | 86 | - | - | - | - | - | - | - | - | - | 94 | - | - | - | - | - |
| STPFaSLSh | - | - | - | 59 | - | - | - | 86 | - | - | - | - | - | - | - | - | - | 70 | - | - | - | - | - | - | 72 | - | - | - | - | - | - | - | - | - | - |
| STPFaSLlo | - | - | - | - | - | - | - | 97 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| TCE | - | - | - | 59 | - | - | - | 88 | - | - | - | - | - | - | - | - | - | - | - | 85 | - | - | - | - | 78 | - | - | - | - | - | - | - | - | - | - |
| TCS | - | - | - | 69 | - | - | - | 79 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 59 | - | - |
| TOAST | - | - | - | - | - | - | - | 83 | - | - | - | - | - | - | - | - | - | - | - | 80 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| TUG-K2.6 | 48 | - | - | - | - | - | - | - | - | 60 | 67 | 60 | - | 43 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| TUG-K3.0-4.0 | - | - | - | 45 | - | - | - | 55 | - | - | - | - | 44 | - | - | - | - | - | - | - | - | - | - | - | 65 | - | - | - | - | 60 | 50 | - | - | - | 45 |
| TUV | 58 | - | - | - | - | - | - | 60 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 62 | - | - | - | - | - |
IV.3 How does the AI’s performance compare to student performance at the undergraduate level?